OpenStudy (anonymous):

It is given that y varies directly as x. If y = 5 when x = 4 determine the constant of variation. A) 4 5 B) 5 4 C) 4 D) 5 @confluxepic

Nnesha (nnesha):

equation for direct variation y = kx where k is constant of variation

OpenStudy (campbell_st):

for direct variation the equation is \[y = k \times x\] k is called the variation constant to find k, substitute y = 5 and x = 4 the solve for k

OpenStudy (anonymous):

1.25

OpenStudy (campbell_st):

thats correct, but you need to write it as a improper fraction

OpenStudy (anonymous):

Sorry I am not good at that at all.

Nnesha (nnesha):

equation is y = kx so replace y and x by their values

OpenStudy (anonymous):

I did that but I can not put it in a fraction

Nnesha (nnesha):

let me know what you get after did how did u substitute x and y values

Nnesha (nnesha):

that**

OpenStudy (anonymous):

1.25

OpenStudy (anonymous):

@Nnesha

Nnesha (nnesha):

thanks okay how did you got 1.25 what did u put into the calculator ??

OpenStudy (anonymous):

I put 5 divided by 4 and 1.25 came up

Nnesha (nnesha):

yes right 5/4 \[\huge\rm 5 = k \times 4\] to solve for k you have to divide both side by 4

Nnesha (nnesha):

\[\huge\rm 1.25 = \frac{ 5 }{ 4 }\]

OpenStudy (anonymous):

Ok so 1.25= k x 1

OpenStudy (anonymous):

So @Nnesha 1.25 is the answer? Of coarse in fraction

OpenStudy (anonymous):

5/4

Nnesha (nnesha):

yes right 5/4 is in fraction form same as 1.25

OpenStudy (anonymous):

Ok thank you!

Nnesha (nnesha):

my pleasure !!

OpenStudy (anonymous):

I have another. I'll tag you if you want to help

Nnesha (nnesha):

i'll try

OpenStudy (anonymous):

ok thank you!